# DeepProtein: Deep Learning Library and Benchmark for Protein Sequence Learning

### v1

@@ 4 Conclusion
Appendix A Hyperparameter Settings
@@
In table 7, we have listed a common settings of hyperparameter used in this library. In terms of learning rate (lr), a higher learning rate which is equal to 0.0001 for graph neural networks would lead to failure in training. For Subcellular and its binary version, a training epoch of 60 is enough for convergence. For small-scale protein datasets such as IEDB [Yi et al., 2018], PDB-Jespersen, and SAbDab-Liberis, a larger learning rate of 0.001 also leads to convergence and the same performance when using CNN, CNN-RNN, and Transformer. For TAP, SAbDab-Chen and CRISPR-Leenay, larger learning rate of 0.0001 is suggested when training graph neural networks.
@@
Appendix B More results of performance on protein sequence learning
In this section, we show three more model performance on Fluorescence, Beta-lactamase and Human PPI dataset.
Figure 4: Performance of selected models including CNN, Transformer, GCN and NeuralFP on Fluorescence, Beta-lactamase and Human PPI. We observe that CNN performs significantly better than GCN on protein function prediction tasks. However, GCN dominates on large-scale PPI datasets such as the Human PPI dataset.
Appendix C More results of empirical memory and runtime complexity
In this section, we show four more empirical complexity results on Fluorescence, Beta-lactamase, SubCellular Binary and Yeast PPI. In Figure 5, the same pattern is found as Figure 3: Transformer took up a large amount of memory to train while the training speed is reasonably fast. Training GNN is slow since the DGL framework computes the message passing on the CPU. CNNs are both fast and memory efficient to train. A different case is found in Figure 6 where for SubCellular Binary and Yeast PPI, GCN is the slowest to train with occupying a large amount of GPU memory space.
Figure 5: Empirical computational complexity on CRISPR and SAbDab Liberis measured by runtime and memory.
Figure 6: Empirical computational complexity on SubCellular Binary and Yeast PPI measured by runtime and memory.
Appendix D Training Curves of DeepProtein
In this section, we plot the metrics (Accuracy or R2) and training loss for SubCellular, PPI Affinity, SAbDab-Chen and TAP.
Figure 7: SubCellular
Figure 8: PPI Affinity
Figure 9: SAbDab_Chen
Figure 10: TAP
Appendix E Few Lines of Codes
Protein Function Prediction
We show an example of CNN on Beta_lactamase with DeepProtein. More case studies can be found in the GitHub repository of DeepProtein. We annotate the code to make researchers better understand the function of code base, which is simpler and clearer than most of the protein learning libraries available.
$>$ import os, sys, argparse, torch, wandb
$>$ from DeepProtein.dataset import *
$>$ import DeepProtein.utils as utils
$>$ import DeepProtein.ProteinPred as models
$>$ path = os.getcwd()
$>$ train_beta = Beta_lactamase(path + ’/DeepProtein/data’, ’train’)
$>$ valid_beta = Beta_lactamase(path + ’/DeepProtein/data’, ’valid’)
$>$ test_beta = Beta_lactamase(path + ’/DeepProtein/data’, ’test’)
$>$ train_protein_processed,train_target,train_protein_idx = collate_fn(train_beta)
$>$ valid_protein_processed,valid_target,valid_protein_idx = collate_fn(valid_beta)
$>$ test_protein_processed,test_target,test_protein_idx = collate_fn(test_beta)
$>$ target_encoding = ’CNN’
$>$ train, _, _ = utils.data_process(X_target=train_protein_processed, y=train_target, target_encoding=target_encoding, split_method=’random’, frac=[0.99998, 1e-5, 1e-5])
$>$ _, val, _ = utils.data_process(X_target=valid_protein_processed, y=valid_target, target_encoding=target_encoding, split_method=’random’, frac=[1e-5, 0.99998, 1e-5])
$>$ _, _, test = utils.data_process(X_target=test_protein_processed, y=test_target, target_encoding=target_encoding,split_method=’random’, frac=[1e-5, 1e-5, 0.99998])
$>$ config = generate_config(target_encoding=target_encoding, \
cls_hidden_dims=[1024, 1024], \
train_epoch=20, \
LR=0.0001, \
batch_size=32, \
$>$ config[’multi’] = False
$>$ torch.manual_seed(args.seed)
$>$ model = models.model_initialize(**config)
$>$ model.train(train, val, test, compute_pos_enc = False)
If we use argparse in Python, this can be easily done in one line, where
python beta.py --target_encoding CNN
--seed 7
--wandb_proj DeepPurposePP
--lr 0.0001
--epochs 100
Note that if you want to observe the result offline without wandb, then set
wandb.init(mode=’offline’, name = "your_project")
inside the training file.
Appendix F Tutorial
Demos can be found under DEMOS folder in DeepProtein.

### v2

@@ 4 Conclusion
Figure 5: Training Loss of selected datasets: PPI_Affinity, SAbDab_Chen, TAP, and SubCellular.
Appendix A Evaluation Metrics
In this section, we describe the basic evaluation metrics for both classification and regression tasks. In the part optimization flow it would be further detailed on the end to end training flow.
Classification metrics. Most classification tasks are binary classification, except subcellular prediction in protein localization prediction, which is a 10-category classification problem, where we use accuracy (acc) (the fraction of correctly predicted/classified samples) as the evaluation metric. In binary classification, there are four kinds of test data points based on their ground truth and the model’s prediction,
  1. 1.
    positive sample and is correctly predicted as positive, known as True Positive (TP);
  2. 2.
    negative samples and is wrongly predicted as positive samples, known as False Positive (FP);
  3. 3.
    negative samples and is correctly predicted as negative samples, known as True Negative (TN);
  4. 4.
    positive samples and is wrongly predicted as negative samples, known as False Negative (FN).
  • Precision. The precision is the performance of a classifier on the samples that are predicted as positive. It is formally defined as $\text{precision}=\textit{TP}/(\textit{TP}+\textit{FP}).$
  • Recall. The recall score measures the performance of the classifier to find all the positive samples. It is formally defined as $\text{recall}=\textit{TP}/(\textit{TP}+\textit{FN})$.
  • PR-AUC (Precision-Recall Area Under Curve). The area under the Precision-Recall curve summarizes the trade-off between the true positive rate and the positive predictive value for a predictive model using different probability thresholds.
  • ROC-AUC Area Under the Receiver Operating Characteristic Curve summarizes the trade-off between the true positive rate and the false positive rate for a predictive model using different probability thresholds. ROC-AUC is also known as the Area Under the Receiver Operating Characteristic curve (AUROC) in some literature.
For all these metrics, the numerical values range from 0 to 1, a higher value represents better performance.
Regression metrics. In the regression task, both ground truth and prediction are continuous values.
  • Mean Squared Error (MSE) measures the average of the squares of the difference between the forecasted value and the actual value. It is defined as $\text{MSE}=\frac{1}{N}\sum^{N}_{i=1}(y_{i}-{\hat{y}}_{i})^{2},$ where $N$ is the size of the test set; $y_{i}$ and $\hat{y}_{i}$ denote the ground truth and predicted score of the $i$-th data sample in the test set, respectively. MSE value ranges from 0 to positive infinity. A lower MSE value indicates better performance.
  • Mean Absolute Error (MAE) measures the absolute value of the difference between the predicted value and the actual value. It is defined as $\text{MAE}=\frac{1}{N}\sum^{N}_{i=1}|y_{i}-{\hat{y}}_{i}|,$ where $N$ is the size of the test set; $y_{i}$ and $\hat{y}_{i}$ denote the ground truth and predicted score of the $i$-th data sample in the test set, respectively. MAE value ranges from 0 to positive infinity. It emphasizes the ranking order of the prediction instead of the absolute value. A lower MAE value indicates better performance.
  • Spearman rank correlation ($\bf{\rho}$), also known as Spearman’s $\rho$, is a nonparametric statistical test that measures the association between two ranked variables. A higher $\rho$ value indicates better performance.
  • R-squared ($R^{2}$) score is defined as the proportion of the variation in the dependent variable that is predictable from the independent variable(s). It is also known as the coefficient of determination in statistics. Higher $R^{2}$ scores indicate better performance.
Appendix B Optimization Flow
Dataset Selection and Processing Flow
As mentioned in the introduction part and Table 1, previous benchmarks either lack 1) the state-of-the-art deep learning methods 2) the diverse real-world data 3) easy-to-use files for researchers outside the computer science domain to use. Hence we collected the data from two main databases which contain approximately 20+ protein tasks which are enough for downstream testing. From them we deleted the tasks that were related to the drug, especially the task drug-target interaction since the DeepPurpose library supported such functionality. For the datasets in the PEER benchmark, DeepProtein just inherited the functions which transformed the data into standard torch datasets, and for the TDC data they were transformed to the standard torch dataset similarly. When loading the dataset, it will load a pair of (protein sequence, aim) or a triple of (protein sequence 1, protein sequence 2, aim), depending on the task type.
Featurization Flow
Since we are talking about training here instead of inference, we ignore the featurization flow of large language models. We mainly consider three types of methods here, which are sequence-based, structure-based and pretraind protein language models.
Sequence-based models take the tokenized SMILES string as the input X,
CNN: $\quad X^{(l)}=X^{(l-1)}*W^{(l)}+b^{(l)}$ (1)
RNN: $\quad h_{t}=\sigma(W_{h}X_{t}+U_{h}h_{t-1}+b_{h})$
Attention: $\quad\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V$
In CNN, $W^{(l)}$ is the weight matrix at layer l which is convoluted by the last hidden layer input $X^{(l-1)}$ and $b^{(l)}$ is the bias. The hidden state is decided by $X^{(l)}$. In RNN, we take each token in $X$ as the input at each time step t, where $W_{h}$ is the weight matrix for the current input token (amino acid) $X_{t}$ and $U_{h}$ is the weight matrix for the last hidden state and $b_{h}$ is the bias. We note that the protein sequence would be long in real-world data, so we truncate them to the maximum length 300, which also avoids memory exploding in RNN. In transformer, for each attention block, we could compute Q, K, V by $W_{Q}X$, $W_{K}X$, and $W_{V}X$, then attention is computed by equation (1). Noted that we could aggregated heads of attention to perform multi-head attention.
Structure-based models take the graph $\mathcal{G}$ as the input, with node features $\mathbf{H}^{(0)}$ and adjacency matrix $\mathbf{A}$. Edge features could be added well if it’s well prepared by the dataset. Especailly for the 2D protein structure, we could only obtain node features by using the features from CNN for instance. GCN, GAT and Graph Transformer’s forms are given by:
GCN: $\quad\mathbf{H}^{(l+1)}=\sigma\left(\tilde{\mathbf{D}}^{-\frac{1}{2}}\tilde{\mathbf{A}}\tilde{\mathbf{D}}^{-\frac{1}{2}}\mathbf{H}^{(l)}\mathbf{W}^{(l)}\right)$ (2)
GAT: $\quad\mathbf{h}_{i}^{(l+1)}=\sigma\left(\sum_{j\in\mathcal{N}(i)\cup\{i\}}\alpha_{ij}^{(l)}\mathbf{W}^{(l)}\mathbf{h}_{j}^{(l)}\right),$
$\quad\alpha_{ij}^{(l)}=\frac{\exp\left(\text{LeakyReLU}\left(\mathbf{a}^{\top}\left[\mathbf{W}^{(l)}\mathbf{h}_{i}^{(l)}\|\mathbf{W}^{(l)}\mathbf{h}_{j}^{(l)}\right]\right)\right)}{\sum_{k\in\mathcal{N}(i)\cup\{i\}}\exp\left(\text{LeakyReLU}\left(\mathbf{a}^{\top}\left[\mathbf{W}^{(l)}\mathbf{h}_{i}^{(l)}\|\mathbf{W}^{(l)}\mathbf{h}_{k}^{(l)}\right]\right)\right)}$
Graph Transformer: $\quad\mathbf{H}^{(l+1)}=\text{Softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d}}+\mathbf{A}\right)\mathbf{V},$
$\quad\mathbf{Q}=\mathbf{H}^{(l)}\mathbf{W}_{Q},\quad\mathbf{K}=\mathbf{H}^{(l)}\mathbf{W}_{K},\quad\mathbf{V}=\mathbf{H}^{(l)}\mathbf{W}_{V}$
MPNN: $\quad\mathbf{m}_{v}^{(l)}=\sum_{u\in\mathcal{N}(v)}M\left(\mathbf{h}_{v}^{(l)},\mathbf{h}_{u}^{(l)},\mathbf{e}_{vu}\right),$
$\quad\mathbf{h}_{v}^{(l+1)}=U\left(\mathbf{h}_{v}^{(l)},\mathbf{m}_{v}^{(l)}\right)$
In GCN, $\mathbf{H}^{(l)}$ is the node representation at layer l, $\tilde{\mathbf{A}}$ is the adjacency matrix with self loop added. $\tilde{\mathbf{D}}$ is the degree matrix corresponding to $\tilde{\mathbf{A}}$. $\mathbf{W}^{(l)}$ is the weight matrix at layer l. in GAT, $\alpha_{ij}^{(l)}$ works as a trainable attention parameter to present the attention betwen node i and node j at layer l. In a general graph transformer, adjacency matrix is added to the attention term which is different from the vanilla self-attention block. Therefore, the complexity is still $O(n^{2})$ if there’re $n^{2}$ nodes in the graph. In the message passing neural network (MPNN), additional edge information $e_{vu}$ is considered for node v and node u.s
For the pre-trained protein languade models (PLM), the general form could be written as:
X=PLM($\textbf{X})$ = PLM ( X ) (3)
where we regard PLM as a white box model. We could get the embedding for the whole protein sequence instead of encoding each amino acid one by one which is more efficient than sequence-based or structure-based encoding.
Training Flow
After we obtained features from the featurizer module, we train the downstream tasks with a linear layer with the weight W and bias $b$. We consider five machine learning task types (not referring to protein learning tasks), which are single protein regression, single protein classification, protein pair regression, protein pair classification, and token (residue) level single protein classification. We introduce them one by one.
Single protein regression task is that given a single protein’s representation X, after applying the linear layer, we got a floating-point number $\hat{y}$, so mean squared error loss between the true value $y_{i}$ and predicted value $\hat{y}_{i}$ is applied during training. For single protein classification task, we apply the softmax function after the linear layer to decide its class. Either a binary cross-entropy loss (BCELoss) or a general cross-entropy loss (CELoss) would be back-propagated during the training:
[…]
$\mathcal{L}_{\text{MSE}}=\frac{1}{N}\sum_{i=1}^{N}(\hat{y}_{i}-y_{i})^{2}$
[…]
$\mathcal{L}_{\text{CE}}=-\sum_{i}y_{i}\log(\hat{y}_{i})$
Protein pair regression is that, the input is a pair of protein ($X_{i},X_{j}$), the aim is to predict its affinity or some other related interaction metrics, labeled $y_{ij}$ here. The representation of two proteins is concatenated before it is applied to a linear layer. The predicted value is $\hat{y}_{ij}$. MSE loss between $\hat{y}_{ij}$ and $y_{ij}$ would be back-propagated. For protein pair classification task, we apply a sigmoid function as all labels are either 0 or 1 in our benchmark. BCELoss is being computed.
[…]
$\mathcal{L}_{\text{MSE}}=\frac{1}{N}\sum_{i,j}(\hat{y}_{ij}-y_{ij})^{2}$
[…]
$\mathcal{L}_{\text{BCE}}=-\sum_{i,j}y_{ij}\log(\hat{y}_{ij})+(1-y_{ij})\log(1-\hat{y}_{ij})$
For residue-level single protein classification, we predict the class for each token (amino acid) for each protein sequence, from the token $X_{1}$ to the the token $X_{T}$ if the length is equal to T. A softmax is applied here after applying a linear layer and CE-loss is calculated. Note that it is computation inefficient to perform residue level prediction when applying graph neural networks and protein language models and could easily reach memory bound so in our benchmark we only tested those datasets with CNN, CNN-RNN and Transformer architectures. $\hat{Y}_{t,c}$ is the probability that $X_{t}$ is being assigned to class c.
[…]
$\mathcal{L}_{\text{CE}}=-\frac{1}{T}\sum_{t=1}^{T}\sum_{c}Y_{t,c}\log(\hat{Y}_{t,c})$
Appendix C Tables of Time and Memory Usage
Table 8: Memory and Time Usage of Different Models
Dataset Model Time (hour) GPU Memory (GB)
Fluorescence CNN 0.01000 0.78000
Fluorescence GCN 0.13200 1.00000
Fluorescence Prot-T5 0.00160 5.20000
Stability CNN 0.01600 0.84480
Stability GCN 0.06100 0.73656
Stability Prot-T5 0.00472 10.19000
Beta CNN 0.00150 0.84700
Beta GCN 0.14000 1.07000
Beta Prot-T5 0.00060 5.20000
Solubility CNN 0.01300 0.84700
Solubility GCN 0.48000 2.70200
Solubility Prot-T5 0.04000 7.33440
SubCellular CNN 0.00600 2.94500
SubCellular GCN 0.07000 8.72600
SubCellular Prot-T5 0.00060 5.71200
SubCellular-Bin CNN 0.00110 1.82600
SubCellular-Bin GCN 0.05200 10.75200
SubCellular-Bin Prot-T5 0.00030 5.62300
PPI_Affinity CNN 0.00400 1.37000
PPI_Affinity GCN 0.03000 1.64600
PPI_Affinity Prot-T5 0.00050 5.16200
Yeast-PPI CNN 0.01000 2.69000
Yeast-PPI GCN 0.32000 12.62000
Yeast-PPI Prot-T5 0.00020 9.04000
Human-PPI CNN 0.14000 0.96000
Human-PPI GCN 1.10000 2.13000
Human-PPI Prot-T5 0.00320 9.38000
TAP CNN 0.01000 1.95000
TAP GCN 0.00300 0.69100
TAP Prot-T5 0.00030 5.16000
SAbDab-Chen CNN 0.00300 0.65000
SAbDab-Chen GCN 0.01000 1.05000
SAbDab-Chen Prot-T5 0.00060 5.16000
CRISPR CNN 0.00020 0.62900
CRISPR GCN 0.00100 0.53800
CRISPR Prot-T5 0.00050 5.16000
Fold CNN 0.01000 0.92200
Fold GCN 0.06000 1.80700
Fold Prot-T5 0.00020 5.22200
Appendix D Prompt Template
Template for ChemLLM-7B:
System Instruction: You are an AI assistant specializing in protein property prediction. Follow the given instruction format. User Prompt Format: <|im_start|>user
What is the {protein_property} of the given protein sequence {protein_sequence}? {instruction}
<|im_end|>
<|im_start|>assistant
Template for LlaSMol-Mistral-7B:
[…]
Instruction and property (task) for each dataset.
Datasets Instruction Property
Fluorescence You should return a floating-point number. Fluorescence intensity
Beta You should return a floating-point number. Increased activity
Stability You should return a floating-point number. Protein stability
Solubility You should return an integer (0 or 1) where 0 is not soluble and 1 is soluble. Protein solubility
Subcellular You should choose an integer within the range [0, 9] to indicate the protein’s location. Location
Subcellular_Binary You should return an integer (0 or 1) where 0 is membrane-bound and 1 is soluble. Location
Tap You should return a floating-point number. Developability
SAbDab_Chen You should return a floating-point number. Developability
CRISPR You should return a floating-point number. Repair outcome
PPI-Affinity You should return a floating-point number. Activity of protein-protein interaction
Yeast-PPI You should return an integer (0 or 1) where 0 is weak and 1 is strong. Activity of protein-protein interaction
Human-PPI You should return an integer (0 or 1) where 0 is weak and 1 is strong. Activity of protein-protein interaction
Fold You should return an integer within the range [0, 1194]. Global structural topology of a protein on the fold level
Secondary You should return an integer within the range [0, 2]. Local structures of protein residues in their natural state
Appendix E Hyperparameter Settings
@@
In table 10, we have listed a common settings of hyperparameter used in this library. In terms of learning rate (lr), a higher learning rate which is equal to 0.0001 for graph neural networks would lead to failure in training. For Subcellular and its binary version, a training epoch of 60 is enough for convergence. For small-scale protein datasets such as IEDB (Yi et al., 2018), PDB-Jespersen, and SAbDab-Liberis, a larger learning rate of 0.001 also leads to convergence and the same performance when using CNN, CNN-RNN, and Transformer. For TAP, SAbDab-Chen and CRISPR-Leenay, larger learning rate of 0.0001 is suggested when training graph neural networks.
@@
Appendix F Competing interests
No competing interest is declared.
Appendix G Author contributions statement
J.X. and T.F. conceived the experiment(s), J.X. conducted the experiment(s), and J.X., Y.Z., and T.F. analyzed the results. J.X., Y.Z., and T.F. wrote and reviewed the manuscript.
Appendix H Acknowledgments
The authors thank the anonymous reviewers for their valuable suggestions. Y.Z. is partly supported by National Science Foundation (NSF) Award No. 2346158.
Appendix I Appendix
You may include other additional sections here.